OpenStudy (anonymous):

Express [4,3] as a linear combination of the given vectors..

OpenStudy (anonymous):

The standad basic vectors i,j

OpenStudy (anonymous):

and \[\left(\begin{matrix}-4 \\ -6\end{matrix}\right)\] \[\left(\begin{matrix}2 \\ 2\end{matrix}\right)\]

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

[4,3] i=[1,0] j=[0,1] 4i=[4,0] 3j=[0,3] 4i+3j=[4,3]

OpenStudy (anonymous):

Hah, why yes. Thank you!

OpenStudy (unklerhaukus):